OpenStudy (anonymous):

Sarah has a total of 52 DVD's and CD's. If the number of CD's is two more than four times the number of DVD's, how many CD's does she have?

OpenStudy (anonymous):

You can write 2 equations with the given information. Let d = the number of DVDs and c = the number of CDs. Can you think of an equation you can write?

OpenStudy (anonymous):

is the awnser 42

OpenStudy (anonymous):

if it is i need another question help with: What is the slope of a line that passes through the point (−2, 3) and is parallel to a line that passes through (3, 7) and (−2, −8)?

OpenStudy (anonymous):

I don't know. I haven't worked it out. How did you get 42?

OpenStudy (anonymous):

10 is the number of dvds and then four times 10: 40 add 2: 42cds + 10dvds = 52 in total

OpenStudy (anonymous):

So you just kinda did it by feel...trying different numbers. If it works go with it. Do you know how to get the slope of the line through 3,7 and -2,-8?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

not really forgot

OpenStudy (anonymous):

Slope = (y2 - y1)/(x2-x1) Sound familiar? m = (-8-7)/(-2-3)=3

OpenStudy (anonymous):

We now know the slope of the line (3) and a point on the line (-2,3). So now use the point-slope line equation: y-y1 = m(x-x1) m is slope (x1,y1) is any point on the line (which we know)

OpenStudy (anonymous):

the slope is three

OpenStudy (anonymous):

Yeah. If were parallel to the line through those 2 points we have to have the same slope as that line.

OpenStudy (anonymous):

Sorrylllgotta go. Family chore thing. bye

OpenStudy (anonymous):

thank you